### v1

# LibriQuote: A Speech Dataset of Fictional Character Utterances for Expressive Zero-Shot Speech Synthesis

Abstract
Text-to-speech (TTS) systems have recently achieved more expressive and natural speech synthesis by scaling to large speech datasets. However, the proportion of expressive speech in such large-scale corpora is often unclear. Besides, existing expressive speech corpora are typically smaller in scale and primarily used for benchmarking TTS systems. In this paper, we introduce the LibriQuote dataset, an English corpus derived from read audiobooks, designed for both fine-tuning and benchmarking expressive zero-shot TTS system. The training dataset includes 12.7K hours of read, non-expressive speech and 5.3K hours of mostly expressive speech drawn from character quotations. Each utterance in the expressive subset is supplemented with the context in which it was written, along with pseudo-labels of speech verbs and adverbs used to describe the quotation (e.g. “he whispered softly”). Additionally, we provide a challenging 7.5 hour test set intended for benchmarking TTS systems: given a neutral reference speech as input, we evaluate system’s ability to synthesize an expressive utterance while preserving reference timbre. We validate qualitatively the test set by showing that it covers a wide range of emotions compared to non-expressive speech, along with various accents. Extensive subjective and objective evaluations show that fine-tuning a baseline TTS system on LibriQuote significantly improves its synthesized speech intelligibility, and that recent systems fail to synthesize speech as expressive and natural as the ground-truth utterances. The dataset and evaluation code are freely available. Audio samples can be found at https://libriquote.github.io/.
Code — https://github.com/deezer/libriquote
Dataset — https://huggingface.co/datasets/gasmichel/LibriQuote
Introduction
Recently, text-to-speech (TTS) systems have significantly improved by scaling to large-scale, multi-domain speech corpora (Shen et al. 2023; Wang et al. 2023; Kharitonov et al. 2023; Jiang et al. 2023a; Anastassiou et al. 2024; Wang et al. 2024; Chen et al. 2024; Wang et al. 2025), exhibiting remarkable speech naturalness, quality and voice following abilities. These large-scale corpora typically contain speech derived from audiobooks (Kahn et al. 2019; Pratap et al. 2020; Kang et al. 2024) or in-the-wild data (He et al. 2024). While in-the-wild speech may include diverse forms of expressivity and speaking style, it is sometimes believed that audiobook speech corpora may lack such expressiveness (He et al. 2024). In contrast, we claim that human reading in audiobooks, especially readings of fictional works, offer extensive prosodic variability, as they often involve an alternation between neutral, non-expressive speech and expressive, sometimes emotional, utterances. Neutral speech usually corresponds in a book to narration parts, these story fragments enunciated by the narrator – may it be a first or third person view, or free indirect speech – while expressive speech is employed during dialogues between fictional characters. Dialogue is an essential narrative function (Genette 1983) that allows fictional character to portray themselves, expressing aspects of their personality, emotional states, and their own representation of the fictional world. As the story unfolds, characters navigate dynamically through their own emotion arc (Vishnubhotla et al. 2024), which is often portrayed during audiobook reading. Indeed, numerous blog posts dedicated to amateurs and written by professional narrators emphasize the importance of impersonation, or the act of bringing life to fictional characters (Gonzalez 2021; Browm 2021; Goodwin 2024).
In this work, we put fictional characters at the forefront, and propose LibriQuote, a large-scale dataset of expressive utterances derived from LibriVox111https://librivox.org recordings. Notably, LibriQuote differs from standard large-scale audiobook datasets (Pratap et al. 2020; Kahn et al. 2019; Kang et al. 2024) by segmenting speech utterances in two ways: character utterances that involve expressive and diverse spech, and narration parts that correspond to a neutral, formal reading style. In addition, we provide as side information the context in which each character utterance appears, supplemented with pseudo-labels of potential speech verbs and adverbs used to describe the utterance (e.g. “he whispered softly”), automatically annotated and further validated by humans. These pseudo-labels provide rich information that we further exploit to construct an expressive data subset. Through an extensive alignment pipeline involving audio preparation and transcription from LibriVox, text matching from Project Gutenberg222https://www.gutenberg.org/ and quotation detection (Bamman, Underwood, and Smith 2014), we collected more than 3 million quotations uttered by fictional characters that we systematically aligned with their book counterpart, resulting in nearly 5400 hours of speech. An overview of the pipeline is provided in Figure 1. Qualitative analysis of the data reveals that LibriQuote contains a large set of emotions and accents, facilitating the training and benchmarking of expressive TTS systems, as depicted in Figure 2.
Our objective and subjective experiments demonstrate the effectiveness of LibriQuote. When evaluated on our test set, modern open-source TTS systems fail to synthesize utterances as expressive and as natural as the ground-truth quotations, suggesting that further research should be conducted to build more expressive systems. When used as a fine-tuning corpora, LibriQuote-train significantly improves speech intelligibility of a baseline TTS system. Besides, we derive an expressive data subset using pseudo-labels of speech verbs and adverbs, which we show is a data-efficient alternative to the full LibriQuote-train set.
Our contribution are as follows:
We release publicly LibriQuote, a dataset derived from audiobooks dedicated to zero-shot TTS. LibriQuote comes with standard train/dev/test splits. LibriQuote-train comprises 5.3K hours of training quotations, as well as 12.7K hours of training narration segments, while LibriQuote-test contains 7.5 hours of quotations distributed among 15 speakers.
We release publicly the code and evaluation samples used in objective and subjective experiments to encourage benchmarking on LibriQuote-test.
We show through extensive objective and subjective evaluations that modern TTS systems fail to synthesize utterances as expressive and as natural as in real audiobooks.
We show that fine-tuning a baseline TTS system on LibriQuote-train significantly improves speech intelligibility on external audiobooks and in-the-wild benchmarks.
Existing Datasets
We briefly review the composition and characteristics of existing datasets used for zero-shot speech synthesis.
Audiobook-based Speech Datasets
Audiobooks have traditionally been the de-facto open-source data available for training TTS systems. The LibriVox project records thousands of audiobooks derived from public-domain books, read by volunteered speakers. LibriSpeech (Panayotov et al. 2015), a 1000-hours corpus of English speech data extracted from LibriVox that covers 2400 speakers, has been extensively used in the literature to train and benchmark Automatic Speech Recognition (ASR) and TTS systems. LibriTTS (Zen et al. 2019) provides an enhanced version of LibriSpeech, with better audio quality and utterance segmentation, but supports a smaller amount of training data, and LibriSpeech-PC (Meister et al. 2023) restored the punctuation and capitalization of LibriSpeech audio transcripts. Libri-Light (Kahn et al. 2019) is the largest English audiobook corpus of open data, with approximately 60,000 hours of English speech derived from LibriVox, but does not provide audio transcriptions, making it unpractical to train supervised ASR or classic TTS systems. The MLS datasest (Pratap et al. 2020) is composed of 50,000 hours of multilingual audiobooks and proposes audio transcriptions that were aligned with the original text. LibriHeavy (Kang et al. 2024) improved Libri-Light by aligning audio segments derived from recordings and their text counterpart in the book from which the segment was taken, but do not provide book contextual information in the vicinity of audio segments. These 30-seconds audio segments are created at sentence boundaries, thus they may contain multiple quotations and narration parts, alternating between neutral (narration) and expressive (quotation) speech. Because of this alternation patterns, complex pitch distributions arise in each training instance, which may hinder the learning of TTS systems. In contrast, we extract speech from a subset of LibriVox, Fiction Novels, separating training instances between neutral (narration) and expressive (quotation) segments, preserving contextual information around quotations. In fiction, quotations from fictional characters is an essential communicative medium between the reader and the characters, revealing their emotions and personality through dialogue.
Expressive Speech Datasets
Audiobook narration typically alternates between neutral, narrative speech and expressive speech during character acting. Székely et al. (2012) design a system to capture expressive speech utterances from audiobooks, and validate their approach on a single Hungarian audiobook. To the best of our knowledge, no work have focused on capturing expressive audio segments in audiobooks at a large scale. However, several efforts towards the creation of expressive read speech corpora have been made. Emotional speech datasets (Busso et al. 2008; Cao et al. 2014; Livingstone and Russo 2018; Zhou et al. 2021) that assign discrete emotion labels to speech utterances have been extensively used to evaluate emotional synthesize of TTS systems (Jiang et al. 2023b; Wang et al. 2024, 2025). The L2-ARCTIC corpus (Zhao et al. 2018) proposes an 11 hour corpus with ten speakers, spanning 5 different accents. EXPRESSO (Nguyen et al. 2023) introduces an high-quality dataset of read and improvised speech across 26 expressive styles. In LibriQuote, we create audio segments that match character utterances and which are thus likely to be more expressive than narration parts. Besides, we extract narrative information indicating how an utterance should be spoken, providing a natural language pseudo-label of expressivity.
LibriQuote Dataset
This section describes the steps involved in preparing LibriQuote, followed by descriptive statistics. Then we present our methodology to extract narrative information for each utterance. Figure 1 provides an overview of the full process.
Dataset Construction
We mainly follow the recipe of LibriHeavy (Kang et al. 2024) to align audio files downloaded from LibriVox to their original text. However, we filter out books that are unlikely to contain fictional characters, such as bibliographical work or other type of non-fictional works. We describe the filtering and alignment stages below.
Audio Preparation
We start by browsing the LibriVox API333https://librivox.org/api/info to collect recording information. We only extract English recordings corresponding to Fiction books. This step is necessary to filter out non-fiction books that are unlikely to contain fictional characters. Each recording is downsampled to 16kHz for further processing, but we provide links to download the original audio files as well, allowing LibriQuote to support systems operating on higher audio quality. Note that LibriVox also contains Dramatic Readings, which are recordings where each fictional character is acted by a unique speaker. We also applied the alignment process described below to these recordings, but do not include them in LibriQuote as they would require an additional speaker diarization step. We plan to release these alignments in future works.
Text Preparation
For each audio recording, we download the corresponding book text from Project Gutenberg. Some books have different text providers, but we found by manual inspection that automatically downloading from these websites was either too complicated, or resulted in texts with very poor quality often due to OCR errors. In such cases, we browse Project Gutenberg to search for the target book, and discard it if not available on the platform. We then perform quotation detection using BookNLP444https://github.com/booknlp/booknlp/blob/main/booknlp/english/litbank˙quote.py on each book. We discard every book containing less than quotations, as we found that these book were typically showing quotation detection errors. For other books, we keep quotation positions in the original text for further alignments.
| Train | Dev | Test | |||
| Count | |||||
| Hours | |||||
| Dur (s) | |||||
| Speakers | |||||
| avg (h) | |||||
| Books | |||||
Audio Transcription
We start by segmenting each audio recording into 30-second segments, with a 2 second overlap at each side. These segments are then transcribed with a Zipformer-Transducer ASR model trained on LibriSpeech555https://github.com/k2-fsa/icefall/pull/1058, and combined by leveraging word-level timestamps to obtain a full timestamped transcript of the recording.
Text-Audio Alignment
Since audio recordings often involve a full chapter, this step is necessary to find the corresponding chapter text within the full original book, and to have fine-grained alignment at the word level. This alignment process involves two stages. In the first stage, close matches between a transcription and the book text are constructed. These close matches correspond to pairs where is an index in the transcription and an index in the book text. Then, a coarse alignment is produced by taking the longest chain of pairs such that and . The second stage produces a final alignment by concatenating Levenshtein alignments (Lcvenshtcin 1966) between the transcription and the text segment produced by the longest chain of pairs. We refer the reader to Kang et al. (2024) for additional information regarding the alignment process.
Quotation Alignment
The output of the text-audio alignment is an array of words in the original text, timestamped in the audio recording. We leverage the timestamps and the output of the quotation detection system to segment each audio into a set of read quotations. Additionally, we build audio segments from narration paragraphs (i.e. parts that are not quotations) in each book. The final construction stage involved assigning contextual information around quotations. Based on the alignments, we match each quotation with a window of around words occurring before and after it, with boundaries defined by paragraphs. With LibriQuote, we release the preprocessed texts as well as full paragraph and quotation alignments. Thus, it is possible to derive a contextual window of any desirable length around quotations.
Descriptive Statistics
We display the main characteristics of LibriQuote in Table 1.
Training Set
The filtering and alignment process resulted in a total of 12723 training hours of narration and 5359 training hours of quotation across 2991 unique books and 3314 unique speakers. As expected, narration parts constitute the majority of audiobook recording in terms of duration, but quotations still represent almost 50% of the total number of training utterances. Indeed, the average audio segment duration is 11.8 seconds for narration parts and 5.5 seconds for quotations. Figure 3 displays the histogram of utterance durations. LibriVox does not provide gender information of its readers, so the balance between male and female speaker in the training set is unclear.
| Verbs | Adverbs | |||||
| Support | 344 | 56 | ||||
| P | R | F | P | R | F | |
| Phi-4 | 0.85 | 0.93 | 0.89 | 0.74 | 0.81 | 0.77 |
| Phi-4 | 0.92 | 0.86 | 0.89 | 0.95 | 0.61 | 0.74 |
Dev & Test Sets
We designed the development and test sets to reflect the following challenge in expressive TTS: given a reference timbre from neutral speech and the text to generate, can we synthesize a quotation with the correct level of expressiveness? Each quotation serves as the expressive speech, and we match them with the nearest narration utterance from the book context that acts as the neutral reference speech. Without the context, the task can be particularly hard as the quotation text itself may be misleading with regard to the true emotion, which may rather be specified in the context. As an example, the exclamation mark in the quotation “It’s so good!” he whispered softly might suggest it shall be uttered with a loud voice, but the narrative information clarifies that it is actually being whispered. Therefore, we also provide contextual information in the vicinity of quotations to encourage further research on predicting expressivity from textual context for audiobook synthesis.
The development dataset contains randomly selected utterances from speakers and books in the train set. It thus does not encompass zero-shot speech synthesis, but aims at evaluating expressive synthesis from neutral speech for seen speakers. The test set contains 8 male and 7 female unseen speakers, for a total of 7.5 quotation hours. We ensure no speaker overlap with LibriSpeech train and dev speakers.
Test Set Analysis
To further analyze LibriQuote test set, we explore its emotional and accent variability. Following Wang et al. (2024), we compute emotion representations using Emotion2Vec666https://huggingface.co/emotion2vec/emotion2vec˙plus˙base (Ma et al. 2024) and accent representations using an ECAPA-TDNN model trained on Common Voice777https://huggingface.co/Jzuluaga/accent-id-commonaccent˙ecapa (Ardila et al. 2020; Zuluaga-Gomez et al. 2023). We display t-SNE projections of emotion representations for LibriQuote-test quotations and their reference narration utterance in Figure 2. For comparison, we also display t-SNE projections of 5734 utterances randomly drawn from all LibriHeavy segments that contain at least one quotation. We see that LibriQuote-test quotations convey a larger set of predicted emotions compared to their narration reference and LibriHeavy segments: only 67% of quotations are predicted as neutral against 87% and 91% for the reference and LibriHeavy respectively. Besides, we found that accent representations are well clustered, indicating a broad coverage of English accents. Further details can be found in the Technical Appendix.
Narrative Information Extraction
| WER | SIM-O | FPC | MCD | A-Sim | E-Sim | CMOS | MOS | |
| Ground Truth | 7.2 | - | - | - | 0.0 | 3.58 | ||
| SparkTTS | 5.5 | 0.45 | 0.65 | 9.1 | 0.61 | 0.69 | -0.97 | 3.26 |
| F5-TTS | 8.1 | 0.52 | 0.63 | 9.5 | 0.66 | 0.70 | -0.98 | 3.47 |
| MaskGCT | 9.3 | 0.55 | 0.63 | 9.5 | 0.65 | 0.72 | -0.98 | 3.36 |
Distinguishing between ”story” (what is happening) and ”discourse” (how it is told) is a key aspect of narrative understanding. NLP research has extensively focused on various aspects of narrative discourse, as theorized by Genette (1983). Recently, a variety of works explored the usage of Large Language Models (LLMs) to model elements of narrative understanding, such as dialogue (Michel et al. 2025), character roles and information extraction (Stammbach, Antoniak, and Ash 2022; Gurung and Lapata 2024), narrativity (Hatzel and Biemann 2024) or structure (Soni et al. 2023).
The current work is grounded in the perspectival issues of fictional characters, including point of view and dialogues, anchored in the voice linking function of Genette’s theoretical framework. In fiction, utterances are sometimes supplemented with narrative information in the form of speech verbs and adverbs (e.g. “he whispered softly”). These indicators have been shown to produce positive effects during reading, such as enhancing reading comprehension (Wolters, Kim, and Szura 2022) or character identification (Van Krieken, Hoeken, and Sanders 2017). For audiobook narrators, they are a valuable resource to set the tone and point of view of the character. Other contextual information such as the emotional context of a character are likely to also impact speech delivery, but are often prone to subjective interpretation and are harder to automatically infer (Vishnubhotla et al. 2024). Thus, by extracting these narrative cues and filtering out less expressive utterances (e.g. “she said”), we can derive a high-quality subset that spans a broad prosodic range.
Method
LLMs have shown promising results in literary dialogue and point of view understanding (Piper and Bagga 2024; Michel et al. 2025). Our experiments involved prompting Phi-4888https://huggingface.co/microsoft/phi-4 to extract narrative information for a target quotation in a contextual window of one paragraph before and after the quotation. We use few-shot prompting with 5 examples. Prompt details can be found in the Technical Appendix. To mitigate extraction errors and keep only informative narrative content, we replace all quotations in context with special markers, such that only narration and structure remains. Besides, we experimented with prompting the LLM to self-report a confidence score on a 1-to-10 scale that we used to prune-out unconfident predictions in order to maximize precision (Phi-4).
Validation
In an initial phase developed to setup guidelines and discussing complex cases, two annotators processed independently 100 quotations. After discussion and once an agreement was reached, a single annotator continued 300 extra instances, resulting in 400 quotations annotated with verbs, adverbs, nouns and adjectives. We provide in Table 2 the results for verbs and adverbs with Phi-4. Results for other LLMs, nouns and adjectives and runtimes can be found in the Technical Appendix. We see that Phi-4 performs relatively well at extracting the narrative information, but still suffers from hallucinations – in particular for adverbs – as shown by the precision score. When pruning out all predictions with a confidence score lower than 10, Phi-4 excels in terms of precision at the cost of lower recall for both verbs and adverbs. A high precision is desirable: we want to avoid quotations to be wrongly affected with expressive speech verbs and adverbs.
Filtering
We extract narrative information on LibriQuote-train using Phi-4, yielding a (potentially empty) set of predicted verbs and adverbs for each quotation. To produce a filtered subset , we first include all quotations that had a non-empty adverb pseudo-label regardless of the predicted speech verb, as adverbs often give precise information of how a quotation is uttered (e.g. “he said cautiously”). Then, based on a manually defined set of expressive speech verbs999We kept 89 out of 201 speech verbs from the following list: https://archiewahwah.wordpress.com/speech-verbs-list/, we include all quotations (which were not kept in the first step) that have a verb falling into that list. As a result, , contains 377,776 quotations (11% of the full quotation train set) for a total of 379 hours. We believe can be used for data-efficiency, or can foster analysis on the alignment between written narrative information and uttered speech.
Experiments
In this section, we describe our main LibriQuote experiments and further evaluations on benchmark datasets.
LibriQuote Test Set
Baselines
To evaluate to which extent state-of-the-art TTS systems are able to synthetize expressive speech, we compare three TTS systems: SparkTTS (Wang et al. 2025), F5-TTS (Chen et al. 2024) and MaskGCT (Wang et al. 2024). These models were chosen based on their availability and performance on standard benchmarks. SparkTTS is an autoregressive TTS model based on Qwen2-0.5B. It generates global and semantic discrete tokens derived from BiCodec (Wang et al. 2025), that are then converted back to raw audio. Global tokens capture speaker information such as timbre, while semantic tokens capture fine-grained semantic information. F5-TTS is a non-autoregressive flow matching TTS system based on Diffusion Transformer (Peebles and Xie 2023a). MaskGCT is a non-autoregressive TTS system, that follows a mask-and-predict paradigm. It follows a two-stage process: semantic tokens derived from a VQ-VAE model with Wav2Vec-Bert 2.0 features are predicted in a first step, and then used to condition the prediction of acoustic tokens. They were all trained with approximately 100K hours of speech data. We use publicly available checkpoints for each model and do not modify generation parameters. Baselines details are available in the Technical Appendix.
| LibriSpeech-PC | SeedTTS test-en | |||
| WER | SIM | WER | SIM | |
| Ground Truth | 2.44 | 0.69 | 2.06 | 0.73 |
| SparkTTS | 3.06 | 0.52 | 2.64 | 0.46 |
| F5-TTS | 2.42 | 0.66 | 1.83 | 0.67 |
| MaskGCT | - | - | 2.62 | 0.72 |
| SparkTTS | 2.10⋆ | 0.51 | 2.07 | 0.42⋆ |
| SparkTTS | 2.00⋆ | 0.51 | 1.90⋆ | 0.42⋆ |
Metrics
Our evaluation follows the cross-sentence design (Le et al. 2023): a narration utterance of 2 to 15 seconds is used as audio context to guide the generation of a quotation utterance. For objective metrics, we report Word Error Rate (WER) computed with Whisper-large-v3 (Radford et al. 2023) to measure speech intelligibility and speaker similarity between the synthesized speech and the original ground-truth speech (SIM-O). We employ a WavLM-large based speaker verification model (Chen et al. 2022) to extract speaker embeddings and calculate cosine similarities. We evaluate prosody similarity with the ground truth quotation using Mel Cepstral Distortion (MCD) and F0 Pearson correlation (FPC) following (Huang et al. 2023). We also evaluate emotion and accent similarity (E-Sim, A-Sim) by leveraging representations from dedicated pretrained models: Emotion2Vec (Ma et al. 2024) and an ECAPA-TDNN model trained on Common Voice (Desplanques, Thienpondt, and Demuynck 2020; Ardila et al. 2020; Zuluaga-Gomez et al. 2023). Similarity scores are calculated by computing the cosine similarity between the ground-truth and synthesized speech representations. For subjective metrics, we employ Mean Opinion Score (MOS, rated from 1 to 5 with 0.5 intervals) to assess naturalness, and Comparative MOS (CMOS, rated from -3 to 3) to measure the degree of expressivity with respect to the ground-truth. We randomly selected 2 samples per speaker in LibriQuote-test, for a total of 30 samples. Each sample was judged by 5 raters. More details can be found in the Technical Appendix.
Analysis
As shown in Table 3, the ground-truth WER is relatively high compared to standard audiobooks datasets such as LibriSpeech. We believe this high WER is likely due to an higher expressiveness in our test samples. SparkTTS achieves the lowest WER (5.5) across models, but the lowest SIM-O. In contrast, MaskGCT achieves the highest WER and SIM-O. F5-TTS and MackGCT synthesize speech with higher accent and emotion similarity than SparkTTS. However, in terms of prosody similarity, SparkTTS performs better, with an higher FPC and lower MCD. The subjective evaluation reveals that all models fail to synthesize speech as expressive as the ground-truth utterances: the latter were consistently preferred over synthesized speech in the CMOS experiment. In addition, ground-truth quotations appear more natural than generated speech, with a MOS of 3.58, while F5-TTS achieves the best MOS of 3.47. These results suggest that the cross-sentence task of LibriQuote-test is challenging, even for modern TTS systems.
Fine-Tuning Experiments
Experiments
We evaluate how and can be used to enhance TTS systems by fine-tuning SparkTTS on these corpus. We chose SparkTTS as a baseline for its simplicity of implementation, which does not requires flow-matching as F5-TTS and MaskGCT.
We compute global and semantic tokens for each training utterances with BiCodec. Then, we fine-tune SparkTTS using the standard language modeling task on semantic tokens with text and global tokens prepended on the two corpora (SparkTTS and SparkTTS). We employ Adam optimizer (Kingma and Ba 2017) with peak learning rate of 1e-5. We train each model for 3 epochs on their respective data. A cosine schedule is used with 10000 warmup steps for SparkTTS and 5000 steps for SparkTTS (note that SparkTTS is only trained for 10% of the total steps of SparkTTS), and we form batches of 32 utterances. During inference, we compute global tokens using reference narration utterances, and prepend the text to synthesize and global tokens before starting the generation.
Datasets & Metrics
We use LibriSpeech-PC (Meister et al. 2023), a punctuation restored version of LibriSpeech – our training data also contains punctuations – with the standard split from Chen et al. (2024) that contains 1127 samples with 4-to-10 seconds audio prompts. We also evaluate on SeedTTS test-en (Anastassiou et al. 2024), which is composed of 1088 samples from Common Voice (Ardila et al. 2020). We do not report scores for the Mandarin variant of SeedTTS as our fine-tuning data only support English speech, although it would be interesting to see how fine-tuning impacts performance on other languages. We report WER computed with Whisper-large-v3 (Radford et al. 2023) and speaker similarity to the reference speech (SIM), computed with WavLM-large (Chen et al. 2022). We compare to SparkTTS, F5-TTS and MaskGCT (only SeedTTS test-en scores are reported for the latter). We report scores computed with samples we generated for SparkTTS on SeedTTS test-en as we were not able to reproduce the original scores.
Analysis
Results are displayed in Table 4. We see that SparkTTS achieves the lowest WER on LibriSpeech-PC, with a large improvement over standard SparkTTS. This gain is also seen on SeedTTS test-en, at the cost of a decrease in SIM. Interestingly, when fine-tuning on the filtered subset , we see a large improvement over SparkTTS, where SparkTTS almost reaches SparkTTS WER and similar SIM, with only 10% of the fine-tuning data and training steps. Thus, reveals as a data-efficient alternative to the full training set for speech intelligibility. These results suggest that training with LibriQuote-train provides substantial improvement to speech intelligibility, while maintaining speaker similarity on audiobook data.
Training from Scratch
| WER | SIM-O | FPC | MCD | |
| SparkTTS | 5.5 | 0.453 | 0.65 | 9.1 |
| SparkTTS | 13.3⋆ | 0.394⋆ | 0.64 | 10.1⋆ |
| context | 11.3⋆ | 0.396⋆ | 0.65 | 9.7⋆ |
| SparkTTS | 6.5 | 0.465⋆ | 0.66⋆ | 8.9⋆ |
| SparkTTS | 10.3⋆ | 0.458⋆ | 0.66 | 9.1 |
Additionaly, we train two variants of SparkTTS from scratch using the quotation subset . We initialize SparkTTS with Qwen2-0.5B weights, and continue training by learning to autoregressively predict the next semantic tokens, conditioned on the text and global tokens. Additionally, we leverage the contextual information around quotation by replacing the text condition with the sequence where and are the text tokens from the left and right context respectively, are the tokens from the text to synthesize, and and are special delimiter tokens. Note that we replace each quotation appearing in and with a special token. Therefore, most of the contextual information can be seen as narrative details. We use a context of one paragraph for both and . We train each variant for 10 epochs with an Adam optimizer, a peak learning rate of 3e-5 and a batch size of 32. A cosine schedule is used with 20000 warmup steps.
Analysis
Results on LibriQuote-test are displayed in Table 5. The model trained from scratch typically suffers from larger WER and lower SIM-O than the original SparkTTS. This is expected, as only 5K hours of training data were used, which is small compared to the 100K hours used by SparkTTS. The same trend follows for FPC and MCD, indicating lower prosody similarity to the ground-truth. A typical problem encountered during generation is word hallucination, which we believe is likely due to the insufficient amount of data to train such autoregressive models. Interestingly, when using contextual information, we see an improvement in WER, FPC and MCD, indicating better speech intelligibility and prosody, which is likely due to the extra narrative details provided. Designing systems capable of leveraging such contextual information might be key to improve the expressiveness of current TTS model.
We also report scores of fine-tuned models in the table: SparkTTS yield a larger WER, but improve slightly the speaker similarity over SparkTTS, while SparkTTS improves over all metric.
Conclusion
This work introduces LibriQuote, a publicly available speech dataset derived from audiobooks, designed for training and benchmarking expressive zero-shot TTS systems. The LibriQuote training set contains 12,720 hours of neutral, non-expressive narration utterances and 5359 hours of more expressive speech in the form of quotations from fictional characters in books, while the test set comprises 7.5 hours of quotations from 15 speakers. Our qualitative analysis reveal that LibriQuote-test covers a large set of emotions and accents. Additionally, we supplement each utterance with the book context in which it was written and provide pseudo-labels for the speech verbs and adverbs used to introduce each quotation. Objective and subjective evaluations show that the TTS systems we evaluated fail to synthesize speech as expressive as the ground-truth utterances and that LibriQuote-train improves speech intelligibility when used as a fine-tuning corpus. We believe that the release of this dataset can foster research on expressive TTS and on contextual conditioning for audiobook synthesis.
Limitations
This work bears several limitations. First, we did not filter LibriQuote utterances that exhibit very high WER between the transcript and ground-truth text as done in previous work. Thus, LibriQuote might contain a small number of outliers, which can hinder learning of ASR or TTS systems. We plan to perform WER-based filtering in a future dataset release. Second, our objective analysis of expressiveness remains limited, with metrics such as FPC and MCD that assume a ground-truth sample is deterministic, hence penalizing diversity in generation. We believe subjective listening tests are more suited for evaluation of expressiveness by providing human ratings. Third, our training experiments were solely conducted with SparkTTS, an autoregressive TTS system which bears its own limitations (such as low speaker similarity), and did not involve other system types. However, we believe that these baseline system results are still a valuable source of information. Finally, we experimented with narrative context conditioning in utterance synthesis, but we believe further research should be conducted on how TTS systems may benefit from it.
Ethical Considerations
This work is purely dedicated to research projects. While TTS models continue bridging the gap towards human naturalness and expressiveness, numerous dangerous applications arise, such as voice spoofing, or the forced replacement of talented voice actors by AI narrators. Therefore, we believe it is of paramount importance to ensure responsible and ethical applications of TTS.
The human study involved participants recruited on a voluntarily basis from our organization, and we ensure each listening test was completed within working hours.
Acknowledgements
This work was performed using HPC resources from GENCI–IDRIS (Grant AD011011668R5)
Appendix A Technical Appendix
Further Descriptive Statistics and Comparisons
In this section, we provide additional information regarding LibriQuote and its comparison to LibriHeavy.
Quotations and Pitch Variation
Audiobook narrators typically alternate between neutral utterances of narration, and expressive utterances in the form of character quotations. Sometimes, they may even completely incarnate a character: for instance, a male speaker might raise his pitch to quote a small female child character. This pitch variation reflects in datasets such as LibriHeavy: audio utterances are typically built by taking around 30 seconds of audio, randomly segmenting at sentence boundaries. Hence, they might contain a various number of quotations, leading to very complex pitch distributions. As a result, Text-to-speech (TTS) systems trained on such audio segments might be hindered by being trained to reproduce these complex distributions without access to meaningful information, such as the fact that a raise in pitch might be favored by the impersonation of a little girl.
To illustrate this effect, we used the small subset of LibriHeavy, and performed quotation detection on each transcription. We found that 25% of audio segments contained 1 to 12 quotations. Then, we calculate pitch standard deviation using Parselmouth (Boersma and Weenink 2021) for each segment and aggregate the results per number of quotations in the segment. The results are displayed in Figure 4. We see that pitch distributions are typically more spread when the number of quotations in the segments increases. This result is supported by looking at Spearman correlations between the number of quotations in the segment and pitch standard deviations (, ). In comparison, LibriQuote-test quotations exhibit a pitch standard deviation of ( bootstrapped 95% confidence interval). Besides, LibriQuote offers 5300 hours of training data distributed among more than 3 millions utterances that are deliberately separated from neutral narration style.
Accent Representations
LibriVox is a collaborative platform where anyone can engage in recording a segment of audiobook. Hence, it contains a varied panel of speakers and accents. We illustrate this in Figure 5 by looking at LibriQuote-test quotations t-SNE projections of accent representations computed with an ECAPA-TDNN model trained on Common Voice101010https://huggingface.co/Jzuluaga/accent-id-commonaccent˙ecapa (Desplanques, Thienpondt, and Demuynck 2020; Ardila et al. 2020; Zuluaga-Gomez et al. 2023). We see a wide range of clusters, that can be interpreted as a large set of accents being represented in the test data.
Emotion Representations
We display in Table 6 the proportion of predicted emotion labels using emotion2vec-plus-base on LibriQuote-test quotations, reference narration and LibriHeavy samples. We used the base variant of emotion2vec plus rather than large as we observed that the large variant was producing collapsed representations, with a very low variance in cosine similarity values. We observe a broader coverage of predicted emotions in quotations compared to their narration references and LibriHeavy audio segments.
Details - Narrative Information
Prompting Experiments
We display in Figure 6 the prompt used to extract narrative information for each quotation. Note that we masked in-context quotations with special markers, and replace the target quotation with “[TARGET]”. This is to ensure that only narrative elements remain in the contextual surroundings of a target quotation.
Additionally, we tested various Large Language Models (LLMs) for the task of narrative information extraction, and also tested the extraction of other elements such as nouns and adjectives. Nouns and adjectives often occur when the utterance description is supplemented by additional information (e.g. “he said with a sigh” for nouns and “he added in a loud, sharp tone” for adjectives). Note that to ensure a high-quality filtering, we tried to maximize the precision of predictions in order to avoid potential hallucinations where some narrative information are spuriously associated to utterances. We see in Table 7 that all tested models seem to perform relatively well on verbs, but some across-model variance is manifested for adverbs. We found that Qwen2.5-72b provides the best overall performance for these two types of information, followed closely by Phi-4. In the end, we decided to use Phi-4 to extract information on the full train set, as it obtained very high precision in 10 less computation time than Qwen2.5-72b. Besides, we found that LLMs dramatically fail at extracting accurately adjectives and nouns, with very low F1 scores. We found during the annotation that extracting such information is harder, even for humans, as introducing adjectives and nouns is less grounded in linguistic rules, and authors might use a more diverse set of linguistic devices to introduce these information. Thus, we decided not to rely on the extraction of such elements.
Overall, extracting verbs and adverbs with Phi-4 (we used vLLM (Kwon et al. 2023) as the inference backend) on the full train set took around 6 hours on a single NVIDIA H100 GPU card.
| LibriQuote (Q) | LibriQuote (R) | LibriHeavy | |
| Neutral | |||
| Angry | |||
| Sad | |||
| Happy | |||
| Surprised | |||
| Fearful | |||
| Disgusted |
Details of Filtering
The result of the extraction is a (potentially empty) set containing extracted verbs and adverbs from Phi-4. To produce the filtered subset , we leverage independently the extracted verbs and adverbs in the following order:
1.
We include all quotations that have a non-empty adverb.
2.
Using the remaining quotations, we include all utterance that have an extracted verb that fall into a predefined list of speech verbs, .
3.
We discard every other quotations.
To build the list of speech verbs, , we started by extracting a large list of 201 potential speech verbs from the web111111https://archiewahwah.wordpress.com/speech-verbs-list/ along with their descriptions. Then, based on verb descriptions, we discarded every verb that might indicate a neutral way of speaking. The resulting verb list contains 89 speech verbs and can be found in Table 8.
| Prompts | Verbs | Adverbs | Adjectives | Nouns | |||||||||
| Support | 400 | 344 | 56 | 28 | 14 | ||||||||
| Runtime (s) | P | R | F1 | P | R | F1 | P | R | F1 | P | R | F1 | |
| Llama3.3-70b | 64 | 0.82 | 0.97 | 0.89 | 0.65 | 0.86 | 0.74 | 0.64 | 0.43 | 0.51 | 0.80 | 0.29 | 0.42 |
| - | 0.91 | 0.91 | 0.91 | 0.8 | 0.54 | 0.65 | 1.0 | 0.05 | 0.09 | 1.0 | 0.07 | 0.13 | |
| Phi4 | 6 | 0.85 | 0.93 | 0.89 | 0.74 | 0.81 | 0.77 | 0.59 | 0.48 | 0.53 | 0.45 | 0.36 | 0.32 |
| - | 0.92 | 0.86 | 0.89 | 0.95 | 0.61 | 0.74 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |
| Qwen3-32b | 19 | 0.86 | 0.93 | 0.89 | 0.69 | 0.92 | 0.79 | 0.5 | 0.76 | 0.6 | 0.43 | 0.71 | 0.54 |
| - | 0.92 | 0.89 | 0.9 | 0.81 | 0.85 | 0.83 | 1.0 | 0.05 | 0.09 | 1.0 | 0.14 | 0.25 | |
| Qwen2.5-72b | 62 | 0.88 | 0.88 | 0.88 | 0.62 | 0.85 | 0.71 | 0.44 | 0.71 | 0.55 | 0.21 | 0.5 | 0.21 |
| - | 0.98 | 0.8 | 0.88 | 0.93 | 0.73 | 0.82 | 0.67 | 0.19 | 0.25 | 0.38 | 0.21 | 0.16 | |
Details - Baselines
In this section, we describe in more details the baselines used in our experiments.
SparkTTS
is an autoregrssive TTS system that builds upon Qwen2.5-0.5M. It also uses BiCodec, a speech codec model that transforms 16kHz sampling rate speech into two types of discrete tokens: global and semantic tokens. Global tokens are derived from representations of an ECAPA-TDNN model fine-tuned for speaker-verification, hence capturing speaker-related information such as timbre. The ECAPA-TDNN representations are mapped to a set of 32 hidden features using learned laten queries, which are then discretized using Finite Scalar Quantization (Mentzer et al. 2023). Semantic tokens are derived from representations of the 11th, 14th and 16th layers of Wav2Vec 2.0 (XLSR-53) (Pasad, Shi, and Livescu 2023). These features are then quantized using single-codebook vector quantization, using factorized codes as done in DAC (Kumar et al. 2023a). Tokens are converted back to raw audio using a Convolutional Neural Network (CNN) Decoder composed of ConvNeXt blocks. The language modeling framework is used to learn to generate semantic tokens, conditioned on the text-to-synthesis tokens and global tokens (which are prepended in the autoregressive sequence).
F5-TTS
is a non-autoregressive diffusion TTS system based on Flow Matching that does not rely on phonetic alignment nor duration predictor, operating on 24kHz sampling rate speech. During training, following the text-guided speech infilling task, speech mel spectrograms are first masked and then concatenated with text tokens. The resulting sequence is inputted to a Diffusion Transformer (DiT) (Peebles and Xie 2023b), which learns to generate the masked spectrogram region from gaussian noise using Flow Matching. During inference, the reference speech mel spectogram serves as the unmasked region, and is concatenated with the text to synthesize the input sequence. Sway sampling is used along with an Ordinary Differential Equation (ODE) solver to generate mel spectrograms that are converted back using Vocos (Siuzdak 2024).
MaskGCT
is also a non-autoregressive TTS system, that follows a two-step masked generative modeling paradigm. It leverages two speech codec models to produce semantic and acoustinc discrete units from 24kHz sampling rate speech. Semantic tokens are derived from representations of the 17th layer of W2v-BERT 2.0, that are converted to discrete units using a single codebook of size 8,192 via classic Vector Quantization (VQ). Acoustic tokens are produced following RVQGAN (Kumar et al. 2023b), and 8 codebooks of size 8,192 are used via Residual VQ (RVQ). To generate raw audio, MaskGCT follows a two-step process: first, it uses a Transformer to generate semantic tokens by iteratively predicting masked tokens conditioned on the text to synthesize, prompt tokens and previously generated semantic tokens ; second acoustic tokens are generated following a similar iterative procedure, but using an architecture similar to SoundStorm (Borsos et al. 2023), and the generative process is conditioned with the prompt acoustic tokens and previously predicted semantic tokens. Finally, predicted acoustic tokens are converted back to raw audio using Vocos. Additionally, MaskGCT employs a Flow-Matching duration predictor to predict the target speech duration during inference.
Details - Subjective Experiments
Human subjects for the listening tests were recruited from our organization via voluntary participation. Participants were indicated that they were part of a study involving the comparison of AI-generated speech and human speech. They were also advised to use headphones and to complete the study in a quiet environment.
For the CMOS experiments, we ask the participants the following question: How expressive is this recording? Please focus on aspects of style (timbre, emotion and prosody), and ignore the aspects of content, grammar, or audio quality. For each instance, we give participants two audios containing one human sample and one generated sample, in a randomized order. We ask the participants to rate if the first audio sounds more expressive than the second on a scale. Screenshot of the CMOS platform can found in Figure 7
For the MOS experiments, we ask the participants the following question: How natural (i.e human-sounding) is this recording? Please focus on examining the audio quality and naturalness, and ignore aspects of style (timbre, emotion, prosody). For each instance, we give participants a single audio, and ask them to rate the naturalness of the sample on an 1-5 Likert scale with 0.5 intervals. Screenshot of the MOS platform can be found in Figure 8
Hardware Information
All experiments done in this paper were conducted on a single compute node equipped with 4 NVIDIA H100 with 80GB of GPU RAM.
Reproducibility Checklist
We would like to address the following points in the Reproducibility Checklist:
All novel datasets introduced in this paper are included in a data appendix. LibriQuote is a speech dataset that include a large amount of audio files, and thus can not be shared straightforwardly in a Data appendix.
For the full dataset release, we plan to release alignments in JSON formats on Huggingface, where a typical LibriQuote instance will comes with start and end anchors (in absolute second position) from an original LibriLight audio file. We will also share scripts to automatically download LibriLight audio recordings and to build the full dataset given the alignment information.
For the data shared in the submission, we provide: 1) a demo page with audio samples used in the listening tests (available at https://libriquote.github.io/) and 2) the alignment information of LibriQuote-test, in case reviewers are interested in looking in details at the test-data.
| Admit | Announce | Argue | Assure |
| Babble | Bark | Bawl | Beg |
| Bellow | Bemoan | Blabber | Bleat |
| Bluster | Boast | Brag | Breathe |
| Cackle | Chant | Cheer | Chirp |
| Chirrup | Cluck | Complain | Confide |
| Cough | Drawl | Exclaim | Falter |
| Fuss | Giggle | Groan | Grumble |
| Growl | Grunt | Hiss | Holler |
| Hoot | Howl | Hum | Implore |
| Jabber | Jibber | Laugh | Moan |
| Mouth | Mumble | Murmur | Mutter |
| Nag | Pant | Pester | Prattle |
| Pronounce | Ramble | Rebuff | Retort |
| Roar | Sass | Scream | Screech |
| Shout | Shriek | Sing | Sigh |
| Snap | Snarl | Snicker | Sniff |
| Snigger | Snivel | Sob | Spit |
| Sputter | Squeak | Squeal | Stutter |
| Taunt | Tease | Trill | Wail |
| Weep | Whimper | Whine | Whisper |
| Whistle | Yell | Yelp | Hesitate |
| Pause |
You are an expert in linguistic. You like to read books and excel at analyzing dialogues in literature.
Given a small narrative passage, where each quote content is masked, your role is to extract speech verbs, adverbs, adjectives and nouns that indicate how a target quotation is being uttered. You will be given a target quotation marked with [TARGET] that occur in the passage. You need to extract speech verbs, adverbs, adjectives and nouns that follow these criteria:
- It must be either a speech-verb, an adverb, a noun or an adjective.
- It must be one word only.
- It must be a speech descriptor of the target quotation.
- If an adverb, it must be a descriptor of one of the speech-verbs describing the target quotation.
- If a verb, ensure that it is a speech-verb and not a verb describing anything else than speech.
Note that multiple speech-verbs can be found and that target quotations can have no associated speech-verbs.
Return a dictionary where keys are the words extracted in the final step and the values is another dictionary with keys ’id’ for the paragraph id of the word, ’type’ for the word type (verb, adverb, adjective, noun) and ’confidence’ an integer between 0 and 10 measuring how confident you are in your prediction, 0 being not confident at all and 10 being sure you are right.
Before creating the dictionary, make sure again that all the criteria above are respected.
Only generate this dictionary and nothing else. Return an empty dictionary if no words were found.
Passage:
[0] Ella handed the notebook to Jay, eyes uncertain.
[1] Jay flipped through the sketches, pausing at one. [QUOTE_1] She nodded.
[2] [TARGET] whispered Ella slowly.
Target quotation: [TARGET]
Answer:
”whispered”: {
”id”: ”2”,
”type”: ”verb”,
”confidence”: 10
”slowly”: {
”id”: ”2”,
”type”: ”adverb”,
”confidence”: 10
Passage:
[0] She went on, half laughing
[1] [TARGET] Then we went to the park, and he said [QUOTE_1]
Target quotation: [TARGET]
Answer:
”went”: {
”id”: ”0”,
”type”: ”verb”,
”confidence”: 9
”laughing”: {
”id”: ”0”,
”type”: ”verb”,
”confidence”: 9
Passage:
[0] I said I had got it on the boat. So then she started for the house, leading me by the hand, and the children tagging after. When we got there she set me down in a split-bottomed chair, and set herself down on a little low stool in front of me, holding both of my hands, and says:
[1] [TARGET]
[2] [QUOTE_1]
Target quotation: [TARGET]
Answer: